5. Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
5x + 5y = 50
10x + 6y = 84
thanks for any help.

Answer by deborabr(181)

5x + 5y = 50
10x + 6y = 84
in the elimination method we delete one of the unknown x or y
is easiest to delete the x in this case
look!
5x + 5y = 50 --> multiply by (-2) to delete x
10x + 6y = 84
-10x-10y=-100 --> sum
10x+6y=84
_____________
-4y=-16 --> change the signal is the same that multiply by (-1) because y is negative
so:
4y=16
y=16/4
y=4
substituting to find x:
5x+5y=50
5x+5*4=50
5x+20=50
5x=50-20
5x=30
x=30/5
x=6
answer: y=4 and x=6